OpenStudy (anonymous):

x^2-8x+8y+32=0 put in y=a(x-h)^2+k or this x=a(y-k)^+h

OpenStudy (anonymous):

what do you want to ask??

OpenStudy (anonymous):

i wanted to put tht first equation into one of the other two..i jus dnt noe which it goes into and how to do it :P

OpenStudy (anonymous):

put tht first equation: x^2-8x+8y+32=0 into y=a(x-h)^2+k or x=a(y-k)^+h form.

OpenStudy (anonymous):

you may want to find out values of h and k.. so convert the equation in the form of first and compare... because power of x is 2 in both equation..

OpenStudy (anonymous):

i jus want to convert the equations into either of those two forms

OpenStudy (anonymous):

\[y= -1/8(x-4)^2 -2\]

OpenStudy (anonymous):

thnk you..i jus need knw how u got tht thou...

OpenStudy (anonymous):

ignore second line

OpenStudy (anonymous):

im kinda confused...